OpenStudy (anonymous):

Pleaseeeeee help!!!:) medal reward! One-sixth of a number is 5 more than the same number. What's the number?

OpenStudy (anonymous):

@some_someone @hedyeh99 They might be able to help cause i sure dont understand this

OpenStudy (anonymous):

1/6 x=5+x

OpenStudy (anonymous):

\[\frac{ x }{ 6 }=x+5\] \[x=6(x+5)\] \[x=6x+30\] \[5x=-30\] \[x=\frac{ -30 }{ 5 }\]

OpenStudy (anonymous):

Yup :D

OpenStudy (anonymous):

Lol, I forgot to write what my equation would equal, it doesn't mater, @some_someone already did that for me, thanks!!!!